There are only two travel agencies, A and B, offering Brazil World Cup 2014 Vacation Packages (flight, hotel accomodation, game tickets and group excursion). The demand fuction for Brazil World Cup 2014 Vacation Packages is given by P=10.000 - Q. The marginal cost is constant and it can be expressed by MC (=ATC)=2.000
a) The travel agencies try to coordinate their actions and set the price and quantity like a single monopolist. Once they set this profit maximizing price and quantity, they plan to split the resulting profil equally. What is the profit for each travel agency if they both adhere to the plan)
b)One of the travel agency, travel agency A, deviates from the plan and sets its price equal to BAM 4.000. What is the profit of travel agency B? (HINT: No one wants to buy overpriced goods!)
c)Both travel agencies set price equal to BAM 4.000, and then split profit equally. What is the profit for each travel agency?.
1
Expert's answer
2017-05-01T10:00:08-0400
P = 10,000 - Q, MC (=ATC) = 2,000, MR = TR' = (P*Q)' = 10,000 - 2Q.
a) If the travel agencies try to coordinate their actions, set the price and quantity like a single monopolist and split the resulting profit equally, then the profit-maximizing quantity is at point where MR = MC, so:
10,000 - 2Q = 2,000,
2Q = 8,000,
Q = 4,000 units,
P = 10,000 - 4,000 = BAM 6,000.
The profit for each travel agency is:
TP = (P - ATC)*Q/2 = (6,000 - 2,000)*4,000/2 = BAM 8,000,000.
b) If one of the travel agency, travel agency A, deviates from the plan and sets its P = BAM 4.000. Then the profit of travel agency B is:
TP = (P - ATC)*Q/2 = (4,000 - 2,000)*4,000/2 = BAM 4,000,000.
c) If both travel agencies set price equal to BAM 4.000, and then split profit equally, then the profit for each travel agency is:
TP = (P - ATC)*Q/2 = (4,000 - 2,000)*4,000/2 = BAM 4,000,000.